1) You comb your hair on a dry day. The comb gained 2.3 x 104 electrons. Each electron has a charge of

-1.6 x 10-19 Coulombs.

a) Is your hair or the plastic comb has a greater affinity for electrons? Why?

b) Assume your hair and comb were all neutral before combing. What is the charge of the comb afterwards?

Eq: (number of electrons gained) x (charge of an electron)

c) What is the charge of your hair afterwards?

Eq: (number more protons than electrons) * (charge of a proton)


1
Expert's answer
2021-02-12T06:10:14-0500

(a) The plastic comb has a greater affinity for electrons, because the comb gained electrons and becomes negatively charged by close contact in combing the hair.

(b) After combing the hair, the comb gained electrons and the charge of the comb becomes negative:


"Q=Ne=2.3\cdot10^4\cdot(-1.6\cdot10^{-19}\ C)=-3.68\cdot10^{-15}\ C."

(c) After combing the hair, the hair has positive charge:


"Q=Ne=2.3\cdot10^4\cdot1.6\cdot10^{-19}\ C=3.68\cdot10^{-15}\ C."